OpenStudy (anonymous):

4 cans of tuna and 2 boxes of rice cost $10. Sic cans of tuna and 2 boxes of rice cost $14. how much does each item cost? Solve by substitution.

OpenStudy (whpalmer4):

Let t be cost of a can of tuna, and r be cost of a box of rice. Then we have: 6 * t + 2 * r = 14, and 4 * t + 2 * r = 10 or 6t + 2r = 14 and 4t + 2r = 10. Conveniently, we've got that 2r in both equations, so rearrange to split that out: 2r = 14 - 6t 2r = 10 - 4t Now set those equal to each other and solve for value of t. 14 - 6t = 10 - 4t 4 = 2t t = 2 Plug in t = 2 into either of the original equations to solve for value of r.